A 58.0 kg box hangs from a rope. What is the tension in the rope if:

Part A

The box is at rest?

Express your answer with the appropriate units.

Part B

The box moves up a steady 4.90 m/s ?

Express your answer with the appropriate units.

Part C

The box has vy = 4.60 m/s and is speeding up at 4.90 m/s2 ? The y axis points upward.

Express your answer with the appropriate units.

Part D

The box has vy = 4.60 m/s and is slowing down at 4.90 m/s2 ?

Express your answer with the appropriate units.
Answer #1

T = tension
m = mass
g = gravitational accel.
a = given accel.

Net F = m*a
a) box at rest, a=0, net F=0, net F = T-mg========> T=mg = 568.4 N

b) box has const. velocity, meaning acceleration still 0. =======> T = 568.4 N

c) T-mg = ma. ========> T=m(g+a) = 58*(9.8 + 4.9) = 852.6 N

d) T = m(g-a) = 58(9.8 -4.9) = 284.2 N